# Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap

### v1

@@ 6 Conclusion
In this paper, we investigated the impact of AR see-through visualizations on human-AI reliance in spatial decision-making under time constraints by comparing them with traditional 2D minimaps. Our user study (N = 32) revealed that AR see-through displays tend to increase over-reliance on AI during spatial decision-making. This effect is primarily attributable to perceptual challenges, such as occlusion and spatial estimation difficulties, as well as AR-specific biases like visual proximity illusions and the influence of high-fidelity visual representations. Nonetheless, AR exhibits unique strengths by enhancing spatial reasoning, as evidenced by significantly reduced pointing errors that suggest improved spatial mapping and navigation. We hope that our findings and design implications will guide future research and the development of solutions that better leverage AR’s advantages to improve human-AI collaboration in high-stakes, spatial contexts.

### v2

@@ 7. Conclusion
In this paper, we investigated the impact of AR embedded visualizations on human-AI reliance in spatial decision-making under time constraints by comparing X-ray with Minimap. Our user study (N = 32) revealed that the X-ray visualization tends to evoke over-reliance on AI during spatial decision-making. This effect may be linked to perceptual challenges (e.g., occlusion and distance estimation) and AR-specific biases, such as visual proximity illusions and the heightened sense of realism and immersion induced by high-fidelity visuals. Nonetheless, the X-ray visualization exhibits unique strengths by enhancing spatial mapping, as evidenced by significantly reduced pointing errors. We hope that our findings and design implications will guide future research and the development of solutions that better leverage AR’s advantages to improve human-AI collaboration in spatial contexts.
Acknowledgements.
The author(s) would like to express sincere gratitude to the anonymous reviewers for their constructive suggestions.